# Some arithmetic aspects of ortho-integral surfaces

### v1

@@ Abstract.
We investigate ortho-integral (OI) hyperbolic surfaces with totally geodesic boundaries, defined by the property that every orthogeodesic (i.e. a geodesic arc meeting the boundary perpendicularly at both endpoints) has an integer cosh-length. We prove that while only finitely many OI surfaces exist for any fixed topology, infinitely many commensurability classes arise as the topology varies. Moreover, we completely classify OI pants and OI one-holed tori, and show that their doubles are arithmetic surfaces of genus 2 derived from quaternion algebras over $\mathbb{Q}$.
@@ 1. Introduction
Since the work of Basmajian, the ortho‑length spectrum has been studied extensively in relation to geometric identities and hyperbolic geometry of the surface. In [Bri11], Bridgeman established an identity connecting the volume of the unit tangent bundle, the ortho‑length spectrum, and the dilogarithm function, with Calegari later offering an alternative proof [Cal11]. Bridgeman and Kahn then generalized this result by showing that the volume of a hyperbolic manifold with nonempty totally geodesic boundary can be recovered from its ortho‑length spectrum [BK10]. In [BT16], Bridgeman and Tan described the common framework underlying these and other geometric identities. For related well-known identities and applications, see [McS98, Mir07, TWZ08, BT14, LT14, DPT23, Par20] and the references therein. Recently, Basmajian, Parlier, and Tan [BPT25] generalized Basmajian’s identity to hyperbolic surfaces with cusps, geodesic boundaries, and cone points by considering prime orthogeodesics in the subsurface obtained by cutting off a portion of the hyperbolic surface using natural collars.
In connection with the hyperbolic geometry of a surface, the ortho-length spectrum has been studied as an analog of the conventional length spectrum. Masai and McShane [MM23] investigated the rigidity and flexibility of the ortho-length spectrum, showing that it uniquely determines the hyperbolic structure on a one-holed torus, constructing examples of non-isometric hyperbolic surfaces with identical ortho-length spectra, and proving that for any fixed hyperbolic surface with non-empty geodesic boundaries, only finitely many non-isometric surfaces share the same ortho-length spectrum. For further recent progress in this direction, see [Que24]. Bavard [Bav05] proved that for a hyperbolic surface $X$ with genus $g$ and $n$ boundary components the length of the shortest orthogeodesic and the total length of the boundary of $X$ cannot both be too large:
@@
The starting point of this article is the work of the first author in [Doa25], where he studied the combinatorial structure of the set of orthogeodesics. In particular, he showed that for a compact hyperbolic surface $X$ with nonempty totally geodesic boundary, the set of orthogeodesics on $X$ can be parametrized by a finite collection of rooted planar trivalent trees which depends on a decomposition of $X$ into right-angled hexagons. Using this tree structure, he provided a recursive method to compute $\mathcal{O}(X)$, thus producing a combinatorial proof of Basmajian’s identity analogous to Bowditch’s combinatorial proof of the McShane identity. When the surface is a pair of pants or a one-holed torus, there always exists a hexagon decomposition consisting of two isometric right-angled hexagons. Due to the symmetry in this decomposition, the cosh-lengths of orthogeodesics on a pair of pants or a one-holed torus can be realized as solutions to a single non-homogenous quadratic equation in three variables.
This connection motivates our investigation into the arithmetic properties of the ortho-length spectrum of hyperbolic surfaces. Since the hyperbolic cosine of half the translation length of a hyperbolic isometry in $\operatorname{PSL}_{2}(\mathbb{R})$ equals half the absolute value of the trace of its matrix, the cosh-length of an orthogeodesic is an analog of the trace of the hyperbolic isometry translating along a closed geodesic. We consider the following definition introduced in [Doa25]:
Definition 1.1.
A hyperbolic surface $X$ is ortho-integral (OI) if all its orthogeodesics have integer cosh-lengths.
In that work, Doan provided examples of OI pants and OI one-holed tori and posed the problem of classifying OI surfaces. In this paper, we present some arithmetic characterization and classification results for these surfaces. Our first main result is a finiteness theorem:
@@
The nonemptiness in 1.2 is established constructively in 5.7 by gluing certain identical OI pants together without twist along their boundaries, either by gluing a pant with itself or with another, to form an OI surface of the desired signature $(g,n)$. The finiteness is proved via a compactness argument in the moduli space of hyperbolic surfaces with nonempty boundary. In fact, we establish this result for a slightly broader class of hyperbolic surfaces (see 3.2 for the full statement). By providing explicit lower bounds for the systole and bounds for the lengths of the boundary components of OI surfaces, we obtain the desired finiteness. A key step in this argument is given by Bavard’s inequality (Equation 2) together with the following integrality property of closed geodesics on OI surfaces:
@@
This integrality extends to a larger class of hyperbolic surfaces (see Remark 3.7). We proved it using the recursive formula in [Doa25] for computing the cosh-lengths of certain orthogeodesics on a pair of pants and an immersion of the pair of pants into the surface. Moreover, through this integrality property and a more careful analysis, we obtain a complete classification of OI pants and OI one-holed tori (see 4.1 and 4.4 for the full statements), confirming that all such surfaces were already found in [Doa25].
Another promising approach to proving 1.2 is to show that the doubles of these surfaces are arithmetic. In fact, a consequence of the volume formula for a maximal order defining an arithmetic Fuchsian group is that there exist only finitely many arithmetic surfaces with bounded area (see, e.g., [Bor81] and [MR03, Chapter 11]). Therefore, if every OI surface doubles to yield an arithmetic surface, finiteness of OI surfaces would follow immediately. Although we did not pursue this strategy directly, we investigated the arithmetic properties of OI pants and OI one-holed tori and verified that the doubles of these OI surfaces are indeed arithmetic surfaces of genus 2 derived from quaternion algebras over $\mathbb{Q}$ (see 5.5, LABEL:tab:OIPantsInvariants and LABEL:tab:OIToriInvariants). It would be interesting to further explore the relationship between arithmetic surfaces and the doubling construction of OI surfaces.
Recall that two hyperbolic surfaces are in the same commensurability class if they share an isometric finite sheeted covering. Motivated by the question of whether there exist infinitely many commensurability classes of pseudomodular surfaces (that is, surfaces of finite area that are not commensurable with the modular surface and that have cusp sets exactly equal to ${}\mathbb{Q}\cup\{\infty\}$; see [LR02] and [LTV18] for the answer to this question), the first author asked in [Doa25] whether there exist infinitely many commensurability classes of OI surfaces. In Section 5.3, we answer this question affirmatively by constructing infinitely many pairwise non-commensurable OI surfaces obtained by gluing together certain OI pants (see Figure 3 for an example).
@@
We note that, unlike the usual arguments for showing that two surfaces are not commensurable by examining their trace fields, which do not work in our case (see Remark 5.12), in the proof of 1.4, we use a completely geometric approach based on comparing the radii of the largest inscribed hyperbolic disks in the universal covers of these surfaces.
@@ 1.1. Organization
In Section 2, we review preliminary results on Mumford’s compactness criterion, the orthospectrum, hyperbolic identities, and the arithmetic invariants of Fuchsian groups. In Section 3, we prove the integrality property of closed geodesics on OI surfaces and establish the finiteness of the set of OI surfaces with a fixed topological type. In Section 4, we classify OI pants and OI one-holed tori. Finally, Section 5 explores connections between OI surfaces and quadratic forms by verifying the arithmeticity of the doubles of OI pants and one-holed tori, proving the nonemptiness of the set of OI surfaces with a fixed topological type, and establishing the infiniteness of commensurability classes of OI surfaces.
@@ 1.2. Acknowledgement
We thank Hugo Parlier, Alan Reid, and Ser Peow Tan for their invaluable discussions and support. We are especially grateful to Greg McShane for his suggestions and for providing the initial data and the idea of 5.2, and to Trong Toan Dao for his assistance with the proof of 4.1. We also thank the organizers of the Computational Aspects of Thin Groups workshop at the Institute for Mathematical Sciences - National University of Singapore for hosting the event where this project began. The first-named author is funded by the Vietnam National Foundation for Science and Technology Development (NAFOSTED) under grant number 101.04-2023.33 and by the Singapore National Research Foundation (NRF) under grant E-146-00-0029-01. The second-named author is funded by the AMS Travel Grant.
@@ 2.2. Moduli space of surfaces, systole, and Mumford’s compactness criterion
The moduli space $\mathcal{M}_{g,n}$ is defined as the space of complete hyperbolic structures with geodesic boundaries on the surface $\Sigma_{g,n}$ (of genus $g$ with $n$ boundary components), considered up to isometries that fix each boundary component individually. To construct $\mathcal{M}_{g,n}$, we begin with the Teichmüller space $\mathcal{T}(\Sigma_{g,n})$, given by
$\mathcal{T}(\Sigma_{g,n})=\left\{(X,f)\,\middle|\,\begin{array}[]{l}X\text{ isa hyperbolic surface with geodesic boundaries},\\ f:\Sigma_{g,n}\rightarrow X\text{ is a homeomorphism}\end{array}\right\}\Big{/}\sim,$
@@ 2.3. Orthospectrum of hyperbolic surfaces
An orthobasis (also known as a hexagon decomposition [BF24] or truncated triangulation [Ush99]) $\mathcal{H}$ of a surface $\Sigma_{g,n}$ (genus $g$ with $n$ geodesic boundary components) is a maximal collection of pairwise disjoint, embedded, essential, simple arcs connecting boundary components, possibly the same boundary. Here an arc is essential if it is not homotopic rel endpoints into one of the boundary component of the surface.
@@
Let =$\mathcal{H}=\{\alpha_{1},\dots,\alpha_{6g-6+3n}\}$ be an orthobasis of a surface $\Sigma_{g,n}$ with genus $g$ and $n$ boundary components. Then the map
@@
Orthogeodesics on $X\in\mathcal{T}(\Sigma_{g,n})$ starting from the same boundary component $\delta$ can be encoded by a rooted planar trivalent tree $T$, known as an orthotree [Doa25, Section 4.1]. The construction of $T$ depends on a choice of an orthobasis $\mathcal{H}$ of $X$ and a choice of a maximal connected open interval $I$ on $\partial X$ disjoint from $\mathcal{H}$. Edges of $T$ are labeled by orthogeodesics in $\mathcal{H}$, and vertices of $T$ are labeled by the hexagon in the decomposition of $X$ by $\mathcal{H}$. We orient all edges away from the root, label the root of $T$ by the initial hexagon that contains $I$ and the edge adjacent to the root of $T$ by the orthogeodesic opposite to $I$ in the initial hexagon. Any finite oriented path starting from the root of $T$ can be thought of as a path on $X$ starting from the initial hexagon crossing a sequence of orthogeodesics in $\mathcal{H}$ transversely. We label the final vertex of the path by the final hexagon in the path on $X$; the left and right edges that point away from this final vertex are labeled by the two remaining left and right orthogeodesics in the final hexagon in the path on $X$.
We observe that each orthogeodesic $\alpha$ starting from $I$ must exit the hexagon containing $I$ through the unique orthogeodesic opposite to $I$. After that, for every hexagon in the decomposition by $\mathcal{H}$ that $\alpha$ enters there are two possibilities. Either $\alpha$ exits the hexagon by crossing one of the two remaining orthogeodesics or $\alpha$ meets the opposite boundary component perpendicularly. Therefore, each finite oriented path from the root corresponds to an orthogeodesic starting at $I\subseteq\delta$, crossing orthogeodesics in $\mathcal{H}$ according to the labels along the path. Thus, orthogeodesics (excluding those in $\mathcal{H}$) starting from $\delta$ correspond to finite paths in $T$. The following theorem connects the cosh-length spectra of orthogeodesics on pants and one-holed tori with group orbits.
@@
We recall some result about short orthogeodesics on $X$. The orthosystole, $\operatorname{osys}(X)$, of a hyperbolic surface $X$ is the length of its shortest orthogeodesic. The orthokissing number, $\operatorname{okiss}(X)$, is the number of orthogeodesics having minimal length. Bavard [Bav05] established a sharp upper bound for the orthosystole of a hyperbolic surface with totally geodesic boundary, restated equivalently by Basmajian and Fanoni [BF24, Theorem 1.1]. The following is an equivalent formulation:

### v2

@@ Abstract.
We investigate ortho-integral (OI) hyperbolic surfaces with totally geodesic boundaries, defined by the property that every orthogeodesic (i.e. a geodesic arc meeting the boundary perpendicularly at both endpoints) has an integer cosh-length. We prove that while only finitely many OI surfaces exist for any fixed topology, infinitely many commensurability classes arise as the topology varies. Moreover, we completely classify OI pairs of pants and OI one-holed tori, and show that their doubles are arithmetic surfaces of genus 2 derived from quaternion algebras over $\mathbb{Q}$.
Key words and phrases:
Hyperbolic surfaces, orthogeodesics, ortho-integral surfaces, arithmetic groups
1991 Mathematics Subject Classification:
57K20, 51B20, 32G15, 15A63, 11F06, 11E20
@@ 1. Introduction
Since the work of Basmajian, the ortho‑length spectrum has been studied extensively in relation to geometric identities and hyperbolic geometry of the surface. In [Bri11], Bridgeman established an identity connecting the volume of the unit tangent bundle, the ortho‑length spectrum, and the dilogarithm function, with Calegari later offering an alternative proof [Cal11]. Bridgeman and Kahn then generalized this result by showing that the volume of a hyperbolic manifold with nonempty totally geodesic boundary can be recovered from its ortho‑length spectrum [BK10]. In [BT16], Bridgeman and Tan described the common framework underlying these and other geometric identities. For related well-known identities and applications, see [McS98, Mir07, TWZ08, BT14, LT14, DPT23, Par20] and the references therein. Recently, Basmajian, Parlier, and Tan [BPT25] generalized Basmajian’s identity to hyperbolic surfaces with cusps, geodesic boundaries, and cone points by considering prime orthogeodesics in the concave core obtained by cutting off natural collars from the surface.
In connection with the hyperbolic geometry of a surface, the ortho-length spectrum has been studied as an analog of the conventional length spectrum. Masai and McShane [MM23] investigated the rigidity and flexibility of the ortho-length spectrum, showing that it uniquely determines the hyperbolic structure on a one-holed torus, constructing examples of non-isometric hyperbolic surfaces with identical ortho-length spectra, and proving that for any fixed hyperbolic surface with non-empty geodesic boundaries, only finitely many non-isometric surfaces share the same ortho-length spectrum. For further recent progress in this direction, see [Que24]. Bavard [Bav05] proved that for a hyperbolic surface $X$ with genus $g$ and $n$ boundary components, the length of the shortest orthogeodesic and the total length of the boundary of $X$ cannot both be too large:
@@
The starting point of this article is the work of the first author in [Doa25], where he studied the combinatorial structure of the set of orthogeodesics. In particular, he showed that for a compact hyperbolic surface $X$ with nonempty totally geodesic boundary, the set of orthogeodesics on $X$ can be parametrized by a finite collection of rooted planar trivalent trees which depends on a decomposition of $X$ into right-angled hexagons. Using this tree structure, he gave a recursive method to compute the set of hyperbolic cosine lengths of all orthogeodesics, with multiplicity, called the ortho cosh-length spectrum:
$\mathcal{O}_{\cosh}(X):=\bigl\{\cosh\bigl(\ell(\mu)\bigr)\;\bigm|\;\mu\text{ is an orthogeodesic on }X\bigr\},$
thus producing a combinatorial proof of Basmajian’s identity analogous to Bowditch’s combinatorial proof of the McShane identity. When the surface is a pair of pants or a one-holed torus, there always exists a hexagon decomposition consisting of two isometric right-angled hexagons. Due to the symmetry in this decomposition, the cosh-lengths of orthogeodesics on a pair of pants or a one-holed torus can be realized as solutions to a single non-homogeneous quadratic equation in three variables.
This connection motivates our investigation into the arithmetic properties of the ortho cosh-length spectrum of hyperbolic surfaces. Since the hyperbolic cosine of half the translation length of a hyperbolic isometry in $\operatorname{PSL}_{2}(\mathbb{R})$ equals half the absolute value of the trace of its matrix, the cosh-length of an orthogeodesic is an analog of the trace of the hyperbolic isometry translating along a closed geodesic. We consider the following definition:
Definition 1.1 ([Doa25]).
A hyperbolic surface $X$ is called ortho integral (OI) if $\mathcal{O}_{\cosh}(X)\subset\mathbb{Z}$, that is, the cosh of the length of every orthogeodesic on $X$ is an integer.
In that work, the first author provided examples of OI pairs of pants and OI one-holed tori and posed the problem of classifying OI surfaces. In this paper, we present some arithmetic characterizations and classification results for these surfaces. Our first main result is a finiteness theorem:
@@
The nonemptiness in 1.2 is established constructively in 5.7 by gluing certain identical OI pairs of pants together without twist along their boundaries, either by gluing a pair of pants with itself or with another, to form an OI surface of the desired signature $(g,n)$. The finiteness is proved via a compactness argument in the moduli space of hyperbolic surfaces with nonempty boundary. In fact, we establish this result for a slightly broader class of hyperbolic surfaces (see 3.2 for the full statement). By providing explicit lower bounds for the systole and bounds for the lengths of the boundary components of OI surfaces, we obtain the desired finiteness. A key step in this argument is given by Bavard’s inequality (Equation 2) together with the following integrality property of closed geodesics on OI surfaces:
@@
This integrality extends to a larger class of hyperbolic surfaces (see Remark 3.7). We proved it using the recursive formula in [Doa25] for computing the cosh-lengths of certain orthogeodesics on a pair of pants and an immersion of the pair of pants into the surface. Moreover, through this integrality property and a more careful analysis, we obtain a complete classification of OI pairs of pants and OI one-holed tori (see 4.1 and 4.4 for the full statements), confirming that all such surfaces were already found in [Doa25].
Another promising approach to proving 1.2 is to show that the doubles of these surfaces are arithmetic. In fact, a consequence of the volume formula for a maximal order defining an arithmetic Fuchsian group is that there exist only finitely many arithmetic surfaces with bounded area (see, e.g., [Bor81] and [MR03, Chapter 11]). Therefore, if every OI surface doubles to yield an arithmetic surface, finiteness of OI surfaces would follow immediately. Although we did not pursue this strategy directly, we investigated the arithmetic properties of OI pairs of pants and OI one-holed tori and verified that the doubles of these OI surfaces are indeed arithmetic surfaces of genus 2 derived from quaternion algebras over $\mathbb{Q}$ (see 5.5, LABEL:tab:OIPantsInvariants and LABEL:tab:OIToriInvariants). It would be interesting to further explore the relationship between arithmetic surfaces and the doubling construction of OI surfaces.
Recall that two hyperbolic surfaces are in the same commensurability class if they share an isometric finite sheeted covering. Motivated by the question of whether there exist infinitely many commensurability classes of pseudomodular surfaces (that is, surfaces of finite area that are not commensurable with the modular surface and that have cusp sets exactly equal to $\mathbb{Q}\cup\{\infty\}$; see [LR02] and [LTV18] for the answer to this question), the first author asked whether there exist infinitely many commensurability classes of OI surfaces ([Doa25]). In Section 5.3, we answer this question affirmatively by constructing infinitely many pairwise non-commensurable OI surfaces obtained by gluing together certain OI pairs of pants (see Figure 4 for an example).
@@
While most non‑commensurability proofs invoke arithmetic or trace‑field invariants (e.g. via quaternion algebras or length‑spectrum comparisons), our argument is entirely geometric. Since the surfaces have nonempty totally geodesic boundary, each gluing of the fixed pair of pants strictly enlarges the convex core and raises its maximal inscribed‑disk radius, producing a strictly increasing sequence of inradii and hence an infinite family of pairwise non‑commensurable surfaces.
@@ 1.1. Organization
In Section 2, we review preliminary results on Mumford’s compactness criterion, the orthospectrum, hyperbolic identities, and the arithmetic invariants of Fuchsian groups. In Section 3, we prove the integrality property of closed geodesics on OI surfaces and establish the finiteness of the set of OI surfaces with a fixed topological type. In Section 4, we classify OI pairs of pants and OI one-holed tori. Finally, Section 5 explores connections between OI surfaces and quadratic forms by verifying the arithmeticity of the doubles of OI pairs of pants and one-holed tori, proving the nonemptiness of the set of OI surfaces with a fixed topological type, and establishing the infiniteness of commensurability classes of OI surfaces.
@@ 1.2. Acknowledgement
We thank Hugo Parlier, Alan Reid, and Ser Peow Tan for their invaluable discussions and support. We are especially grateful to Greg McShane for his suggestions and for providing the initial data and the idea of 5.2, and to Trong Toan Dao for his assistance with the proof of 4.1. We thank the anonymous referee for several valuable comments and suggestions that have improved the paper. We also thank the organizers of the Computational Aspects of Thin Groups workshop at the Institute for Mathematical Sciences - National University of Singapore, for hosting the event where this project began. The first-named author is funded by the Vietnam National Foundation for Science and Technology Development (NAFOSTED) under grant number 101.04-2023.33 and by the Singapore National Research Foundation (NRF) under grant E-146-00-0029-01. The second-named author is funded by the AMS Travel Grant.
@@ 2.2. Moduli space of surfaces, systole, and Mumford’s compactness criterion
The moduli space $\mathcal{M}_{g,n}$ is defined as the space of complete hyperbolic structures with geodesic boundaries on the surface $\Sigma_{g,n}$, considered up to isometries that fix each boundary component individually. To construct $\mathcal{M}_{g,n}$, we begin with the Teichmüller space $\mathcal{T}(\Sigma_{g,n})$, given by
[…]
@@ 2.3. Orthospectrum of hyperbolic surfaces
An orthobasis (also known as a hexagon decomposition [BF24] or truncated triangulation [Ush99]) $\mathcal{H}$ of a surface $\Sigma_{g,n}$ is a maximal collection of pairwise disjoint, embedded, essential, simple arcs connecting boundary components, possibly the same boundary component. Here an arc is essential if it is not homotopic rel endpoints into one of the boundary components of the surface.
@@
Let $\mathcal{H}=\{\alpha_{1},\dots,\alpha_{6g-6+3n}\}$ be an orthobasis of a surface $\Sigma_{g,n}$. Then the map
@@
Orthogeodesics on $X\in\mathcal{T}(\Sigma_{g,n})$ starting from the same boundary component $\delta$ can be encoded by a rooted planar trivalent tree $T$, known as an orthotree [Doa25, Section 4.1]. The construction of $T$ depends on a choice of an orthobasis $\mathcal{H}$ of $X$ and a choice of a maximal connected open interval $I$ on $\partial X$ disjoint from $\mathcal{H}$. Edges of $T$ are labeled by orthogeodesics in $\mathcal{H}$, and vertices of $T$ are labeled by the hexagons in the decomposition of $X$ by $\mathcal{H}$. We orient all edges away from the root, label the root of $T$ by the initial hexagon that contains $I$ and the edge adjacent to the root of $T$ by the orthogeodesic opposite to $I$ in the initial hexagon. Any finite oriented path starting from the root of $T$ can be thought of as a path on $X$ starting from the initial hexagon crossing a sequence of orthogeodesics in $\mathcal{H}$ transversely. We label the final vertex of the path by the final hexagon in the path on $X$; the left and right edges that point away from this final vertex are labeled by the two remaining left and right orthogeodesics in the final hexagon in the path on $X$. See Figure 1 for an example of an orthotree of a pair of pants.
Figure 1. Consider a pair of pants with a standard orthobasis with labels $\{P,B,R\}$ for pink, blue and red. The top four figures show four orthogeodesic arcs with cutting sequence $P$, PRPR, PRPRPRPR and PRBRPRBR starting from an interval II. The bottom figures show the orthotree parametrizing orthogeodesic starting from $I$.
We observe that each orthogeodesic $\alpha$ starting from $I$ must exit the hexagon containing $I$ through the unique orthogeodesic opposite to $I$. After that, for every hexagon in the decomposition by $\mathcal{H}$ that $\alpha$ enters, there are two possibilities. Either $\alpha$ exits the hexagon by crossing one of the two remaining orthogeodesics or $\alpha$ meets the opposite boundary component perpendicularly. Therefore, each finite oriented path from the root corresponds to an orthogeodesic starting at $I\subseteq\delta$, crossing orthogeodesics in $\mathcal{H}$ according to the labels along the path. Thus, orthogeodesics (excluding those in $\mathcal{H}$) starting from $\delta$ correspond to finite paths in $T$. The following theorem connects the cosh-length spectra of orthogeodesics on pairs of pants and one-holed tori with group orbits.
@@
We recall some result about short orthogeodesics on $X$. The orthosystole, $\operatorname{osys}(X)$, of a hyperbolic surface $X$ is the length of its shortest orthogeodesic. The orthokissing number, $\operatorname{okiss}(X)$, is the number of orthogeodesics of length $\operatorname{osys}(X)$. Bavard [Bav05] established a sharp upper bound for the orthosystole of a hyperbolic surface with totally geodesic boundary, restated equivalently by Basmajian and Fanoni [BF24, Theorem 1.1]. The following is an equivalent formulation: